Inc. is concerned about the taxes paid by the company in 2018. In addition to $4.3 million of taxable income, the firnm received $304,000 of interest on state-issued bonds and $140,000 of dividends on common stock it owns in Oakdale Fashions, Inc (Use Iable23) Calculate Hunt Taxidermys tax liability (Enter your answer in dollars not in millions. Round your answer to the nearest dollar amount.) Calculate Hunt Taxidermys average tax rate. (Round your answer to 2 decimal places.) tax rate
table 2.3 Corporate Tax Rates as of 2015 Pay this Amount on Base Income Plus this Percentage on Anything Over the Base Taxable Income S0-$50,000 50,001-$75,000 75,001-$100,000 100,001-$335,000 335,001-$10,000,000 10,000,001-$15,000,000 15,000,001-$18,333,333 Over $18,333,333 7,500 3,750 22,250 113,900 ,400,000 5,150,000 6,416,667 1596 25 34 39 34 35 38 35
Answer #1

1. computation of Hunt Taxidermy inc tax liability:

a. interest on state issued bonds are exempted from taxes

b. dividend will be taxed to 30%

so,

Income given $4,300,000

dividend income $140,000 X 30%= $42,000

Taxable income= 4300000+42000= $4,342,000

Tax Liability from the rates given:

113,900+ 34%(4,342,000-335001)= 113,900+1,362,380 = $1,476,280

2. Average tax rate:

Total tax/ total taxable income*100 = 1476280/4342000*100 = 34%